OpenStudy (anonymous):

How do you solve: -2(t-7)-(t+5)=5 ..?

OpenStudy (anonymous):

solve for t ?

OpenStudy (anonymous):

Or simplify ?

OpenStudy (anonymous):

t= 4/3

OpenStudy (anonymous):

Can you help me break down how you did that?

OpenStudy (anonymous):

-2(t-7)-(t+5)=5 simplify the equation first. MDAS rule applies here.. for -2(t-7): -2t+14 for -(t+5): -t+5 combine the two equations.. -2t+14-t+5=5 now you can solve for t by combining in like terms. -3t+19=5 add both sides by -19 to eliminate the +19.. -3t=5-19 -3t=-14 divide both sides by -3 to eliminate -3 \[t = -14/-3\]